Tohban Report 2014-01-22

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 08 Jan 2014
End Date: 15 Jan 2014
Tohban: Nicole Duncan
Tohban email:
Next Tohban: Mitsuo Oka
List all reports


Solar Activity

Solar activity was very low this week, with zero M class flares, and 20 C class flares. There are currently 11 ARs on disk, only one of which has shown activity in the past day. This spot is in the Eastern hemisphere, and has a chance of some additional C-class flares in the next few days. Stereo B shows 2 ARs that will rotate onto the disk later in the week, with no additional spots in the coming few days.

How many GOES flares occurred?

 Flares above B, C, M, X class were      1    20     0     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      0    4     0     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      0     9     1     0

There were RHESSI flares/GOES flares 197 / 32 over the time range 08-Jan-14 15-Jan-14

Memory Management

Maximum fill of 30%. The SSR is not emptying at the end of the pass set, but gets below the 10% level. This is likely because of skipped passes in the past few days.

Spacecraft Status

The cold plates 1 and 2 are at 119.1 K and 117.6 K, respectively. This is up about .5 degree from last week. Cryocooler is nominal. Power is nominal. Due for spin-up this week, thursday or friday.

Data Gaps

GAPS IN APP_ID = 102 (VC3-MONITOR RATES) WITH PACKET RATE LT 30 GAP START TIME GAP END TIME GAP (SEC) 2014-01-21T09:20:00.000 -- 2014-01-21T09:25:00.000 300.00000 - replayed and filled in 2014-01-17T17:45:00.000 -- 2014-01-17T17:50:00.000 300.00000 - This is the bad App ID packet. This is usually seen about once/ month.

Detector issues

During the last meeting we decided to allow D9 to evolve naturally, avoiding commanding while it remains within a reasonable condition. Over the past week, the rear fast LLD slowly increased from a rate of 1,000 to ~ 5,000. The livetime has decreased as a result of the increased rate, dipping down to ~95% during times of high background. The D9 average lifetime is hovering around 97%. No action has been taken. In the next few days, its projected that the rate will reach 10,000 and at that time the rear fast LLD threshold will likely need to be increased.

There were numerous spikes which appeared across all (or many) detectors. Its thought that these spikes are from collisions, or other space-craft wide issues, not detector arching.

Other notes

The team is gathering data to determine if we should do another anneal. Going to look at the D4 response during the unsegmented ~year in 2012.

Spacecraft Management

Decimation Normal/Vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) Enabled, during entire eclipse
Night time data (rears) Enabled, during entire eclipse
Require extra passes? No?
Requirement for moving pointer? No
Attenuator operation Normal
Detector problems? Detector 9 rear fast rates continue to increase.
Personal tools